OpenStudy (anonymous):

Two angles of a triangle are 47 and 31, find the third angle? How did you come up with the answer 102?

myininaya (myininaya):

3 angles in a triangle sum of 3 angles is 180 180-47-31=102

OpenStudy (anonymous):

All angles of a triangle should add up to 180 degrees. 180-47-31 = 102

OpenStudy (katrinakaif):

Traignles angles add up to 180 degree. The two angles are added up to equal 78. We Substract from 180 - 78 = 102